Whitney'S Appliances is having a sale on 13 TV sets. They are displayed in order of increasing price from left to right. The price of each set differs by $20 from either set next to it. For the price of the set at the extreme right, a customer can buy both the second and seventh sets. What is the price of the least expensive set?
Answer by rapaljer(4551)

Let x = price of TV on the left side (least expensive)
x+20 = 2nd TV
x+40 = 3rd TV
x+60 = 4th TV
x+80 = 5th TV
x+100 = 6th TV
x+ 120 = 7th TV
and so onm notice the pattern is x + 20(n-1)
x + 12*20 = price of 13th TV on the right side (most expensive)
x + 240 = 13th TV

Price of the TV on the right = Price of 2nd + 7th TV
x+240 = x+20 + x+120
x+240 = 2x + 140

Subtract x from both sides:
x+240 - x = 2x -x + 140
240 = x + 140
100 = x

Least expensive TV = $100.

Check: Most expensive TV = $100 + 12*20 = $340
2nd TV = $120
7th TV = $220

GOOD PROBLEM!!

R^2 at SCC